A man standing on the platform of astation , records the frequency of whistles of two trains , one train entering and other train living the station with equal velocities of 4 m/s how many beats are heard by the man persecond if the frequency of each whistle is 240 hertz and the velocity of sound is 320 m/s ?
 

Grade:12

1 Answers

ANMOL SETH
6 years ago
n(of train entering)=240*(320/320-4) =243Hzn(of train leaving) =240*(320/320+4) =237No.of beats heard(Δn) = 243-237=6Therefore 6 beats will be heard by the observer.Thanks